what is the break-even function
c(x)=17x & R(x)=540-10x

Answer by stanbon(75887)

what is the break-even function
c(x)=17x & R(x)=540-10x
---------------
C(x) is Cost
R(x) is Revenue
----
When C(x) = R(x) you break-even.
------
Breal-evem Eqiation: 17x = 540-10x
25x -540 = 0
=================
Cheers,
Stan H.